OpenStudy (anonymous):

if x+y=7 and y+z=8 and x+z=4 then x+y+z=what

OpenStudy (anonymous):

x=7-y z=8-y therefore, (7-y)+(8-y)=4 => y=11/2 hence, x=7-11/2 and z=8-11/2 then just sum them all.

OpenStudy (anonymous):

setting up a system of equations (or matrix) and doing row reduction gives x = 3/2, y = 11/2, and z = 5/2.

OpenStudy (anonymous):

x+y+y+z+z+x=8+7+4 2(x+y+z)=19 x+y+z=19/2